OpenStudy (s3a):

Partial fractions problem I am having: How do I decompose (2 + x^4) / [x(1 + x^4)] into 2/x - x^3/(1 + x^4) (which is the answer - but my book skips the partial fractions steps).

myininaya (myininaya):

\[\frac{2+x^4}{x(1+x^4)}=\frac{A}{x}+\frac{Bx^3+Cx^2+Dx+E}{1+x^4}\] I believe this is how you do do it.

myininaya (myininaya):

now you find A,B,C,D, and E.

OpenStudy (s3a):

I was actually expected to compute all that; I assumed that I must have been wrong, because it seemed too tedious!

OpenStudy (s3a):

Thank you.

myininaya (myininaya):

It actually isn't it hard to compute.